OpenStudy (anonymous):

Vectors : if we have Ax=2cm and Ay=5cm and Bx=4cm and By=-6cm if A-B+2C=0 WHAT'S THE VALUES OF Cx AND Cy is it Cx=1cm and Cy=5.5cm?

OpenStudy (amoodarya):

if A-B+2C=0 (ax,ay)-(bx,by)+2(cx,cy) =0 put them in this and find cx ,cy easily

OpenStudy (anonymous):

So it's like I said ? since ax-bx=-2 and ay-by=11

OpenStudy (anonymous):

so it's cx=1 and cy=-5.5

OpenStudy (amoodarya):

yes ,sure

OpenStudy (anonymous):

thank u